40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 335 of 335  Not logged in ELOG logo
ID Date Author Type Category Subject
  25   Mon Oct 29 11:07:22 2007 waldmanSoftware InstallationOMCSoftware install on OMS
[Alex, Sam]

We spent a little time this morning working on OMS and getting things restarted. A few changes were made. 1) We put openmotif on OMS so that the burtrb doesn't throw that crappy libXm any more. 2) We upgraded OMS to a 32 kHz sampling rate from 2 kHz. All the filters will have to be changed. We also added a PDH filter path to maybe feedback PDH signals cuz that will be cool. Maybe someday I will write up the very cool channel adding procedure.
  24   Mon Oct 29 09:46:50 2007 steveRoutineVACvac & pem trend
Pumpdown 64 pumped by maglev for 125 days
pd64-m-d125

Rob, can you tell me, when did the fire start on this plot?
Attachment 1: pd64md125.jpg
pd64md125.jpg
  23   Mon Oct 29 09:16:31 2007 steveRoutineVACthe rga is back
We had no filament current since last power glitch of Oct. 8, 2007
First I thought that the filament was lost, but it was only bad contact.
The rga head pins were oxidized. Rga was turned back on last Friday.
It's temp is 55.3C normal
Attachment 1: rgaisback.jpg
rgaisback.jpg
  22   Sun Oct 28 03:03:42 2007 ranaConfigurationIOOThree Way Excitement
We've been trying to measure the MC mirror internal mode frequencies so that we can measure
their absorption before and after drag wiping.


It looked nearly impossible to see these modes as driven by their thermal excitation level;
we're looking at the "MC_F" or 'servo' output directly on the MC servo board.

Today, I set up a band limited noise drive into the 'Fast POS' inputs of the 3 MC coil
driver boards (turns out you can do this with either the old HP or the SR785).

Frequencies:
MC1     28.21625 kHz
MC2     28.036   kHz
MC3     28.21637 kHz

I don't really have this kind of absolute accuracy. These are just numbers read off of the SR785.

The other side of the setup is that the same "MC_F" signal is going into the SR830 Lock-In which
is set to 'lock-in' at 27.8 kHz. The resulting demodulated 'R" signal (magnitude) is going into
our MC_AO channel (110B ADC).

As you can see from the above table, MC1 and MC3 are astonishingly and annoyingly very close in
frequency. I identified mirrors with peaks by driving one at a time and measuring on the spectrum
analyzer. I repeated it several times to make sure I wasn't fooling myself; it seems like they
are really very close
but distinct peaks. I really wish we had chipped one of these mirrors
before installing them.



Because of the closeness of these drumhead modes, we will have to measure the absorption by making long
measurements of this channel.
  21   Sat Oct 27 19:00:44 2007 waldmanConfigurationOMCHanging, locked OMC with REFL extracted.
I got the OMC locked to the fiber output today. It was much more difficult than I expected and I spent about 30 minutes or so flailing before stopping to think. The basic problem is that the initial alignment is a search in 4-dimensional space and there is naturally only one signal, the reflected DC level, to guide the alignment. I tried to eyeball the alignment using the IR card and "centering" the beams on mirrors, but I couldn't get close enough to get any light through. I also tried to put a camera on the high reflector transmission, but with 1.5 mW incident on the cavity, there is only 1.5 microwatts leaking through in the best case scenario, and much, much less during alignment.

I resolved the problem by placing a high reflector on a 3.5 inch tall fixed mount and picking off the OMC transmitted beam before it reaches the DC diodes. I took the pickoff beam to a camera. The alignment still sucked because even though the beam cleanly transmitted the output coupler, it wasn't anywhere close to getting through the OTAS. To resolve this problem, I visually looked through the back of M2 at M1 and used the IR card to align the beam to the centers of each mirror. That was close enough to get me fringes and align the camera. With the camera aligned, the rest was very easy.

I restored the PDH setup we know and love from the construction days and locked the laser to the OMC with no difficulty. The laser is in Rana's lab so I send the +/- 10V control signal from the SR560 down a cable to 058E where it goes into the Battery+resistor box, the Throlabs HV amplifier, and finally the FAST channel of the NPRO. BTW, a simple experiment sows that about 35 +/- 3 V are required to get an FSR out of the NPRO, hence the Thorlabs HV. The EOM, mixer, splitter, etc is on the edge of the table.

With this specific OMC alignment, ie. the particular sitting on EQ stops, it looks like all of the ghost beams have a good chance of coming clear. I can fit a 2 inch optic in a fixed mount in between the end of the breadboard and the leg of the support structure. A picture might or might not be included someday. One of the ghost beams craters directly into the EQ stop vertical member. The other ghost barely misses M2 on its way down the length of the board. In its current configuration, the many REFL beam misses the leg by about 1.5 inches.
  20   Fri Oct 26 21:48:40 2007 waldmanConfigurationOMCFiber to 056
I set up a 700 mW NPRO in Rana's lab and launched it onto a 50m fiber. I got a few mW onto the fiber, enough to see with a card before disabling the laser. The fiber now runs along the hallway and terminates in rm 056. Its taped down everywhere someone might trip on it, but don't go out of your way to trip on it or pull on it because you are curious. Tomorrow I will co-run a BNC cable and attenuate the NPRO output so it can only send a few mW and so be laser safe. Then we can try to develop a procedure to align the beam to a suspended OMC and lock our suspended cavity goodness.

Notes to self: items needed from the 40m
  • ND10 and ND20 neutral density filter
  • EOM and mount set for 4 inch beam height
  • Post for fiber launch to get to 4 inch
  • Mode matching lens at 4in
  • 3x steering mirror at 4in
  • RF photodiode at 4in
  • Post for camera to 4in
  • Light sheild for camera
  • Long BNC cable
Some of these exist at 056 already
  19   Fri Oct 26 17:34:43 2007 waldmanOtherOMCOMC + earthquake stops

[Chub, chris, Pinkesh, Sam]

Last night we hugn the OMC for the first time and came up with a bunch of pictures and some problems. Today we address some of the problems and, of course, make new problems. We replaced the flat slotted disks with the fitted slotted disks that are made to fit into the counterbore of the breadboard. This changed the balance slightly and required a more symmetric distribution of mass. It probably did not change the total mass very much. We did find that the amount of cable hanging down strongly affected the breadboard balance and may also have contributed to the changing balance.

We also attached earthquake stops and ran into a few problems:

  • The bottom plate of the EQ stops is too thick so that it bumps into the tombstones
  • The vertical member on the "waist" EQ stops is too close to the breadboard, possibly interfering with the REFL beam
  • The "waist" EQ stops are made from a thin plate that doesn't have enough thickness to mount helicoils in
  • Helicoil weren't loaded in the correct bottom EQ stops
  • The DCPD cable loops over the end EQ stop looking nasty but not actually making contact

However, with a little bit of jimmying, the EQ stops are arrayed at all points within a few mm of the breadboard. Meanwhile, Chub has cabled up all the satellite modules and DCPD modules and Pinkesh is working on getting data into the digital system so we can start playing games. Tonight, I intend to mount a laser in Rana's lab and fiber couple a beam into the 056 room so we can start testing the suspended OMC.
  18   Fri Oct 26 16:19:29 2007 Tobin FrickeRoutineIOOMC resonances
We would like to measure the absorption of the mode cleaner optics. The plan is to repeat <a href="http://ilog.ligo-wa.caltech.edu:7285/mLIGO/Cleaning_the_Mode_Cleaner">Valera's experiment</a> in which we track the MC's thermal resonances to infer their power absorption. Last night Rana and I hooked up a lock-in amplifier to heterodyne the MC servo signal by 28 kHz and piped the output into an ADC using the MC_AO channel. We did not find any resonances.

Valera recommends we drive the POS of the three MC optics with bandlimited noise to excite the resonances.
  17   Fri Oct 26 09:10:17 2007 steveRoutinePEMPEM &PSL trend
The fires are out, lab particle counts are up.
Psl HEPAs are at 100% and mobel HEPAs are just turned on
20 days plot and 5 hrs plot below
Attachment 1: counts&psl.jpg
counts&psl.jpg
Attachment 2: 5dcounts.jpg
5dcounts.jpg
  16   Thu Oct 25 23:35:36 2007 waldmanOtherOMCHang the OMC!
[Pinkesh, Sam]

We tried, convicted and hung the OMC today. The OMC was found guilty of being overweight, and unsymmetrically balanced. The unsymmetry was kind of expected and was corrected with a hefty stack of counterweights positioned over the counterweighting holes. The stacks will be measured at some future date and correctly sized objects machined. The overweightness showed up when the level hanging breadboard was about 5 mm low. This showed up in the board height above the table as well as the OSEM flag positions within their holes. The problem was remedied with a liposuction of the intermediate mass. We removed both small vertical cylinder weights that Chris added, and then we removed the heavy steel transverse weight that can be used to adjust the tip around the long axis (I forgot what its called).

The top of the breadboard ended up about 154 mm off the table. The breadboard is 39 mm thick, and the optics are centered (30 - 12.7) = 17.3 mm below the surface for a as hanging beams height of 154 -39 - 17.3 = 97.7 mm or about an 0.150 inches lower than we were aiming for. Can I get a refund?

We screwed up in multiple ways:
  • The slotted disks that capture the wires do not have the alignment bore used to center the wire in the hole
  • We didn't correctly route the far field QPD cable so it runs funny
  • We didn't have a tool which could be used to get two of the DCPD preamp box mounting screws (which are M3's chub!)
  • We don't have the cable clamps to tie off the electrical cables to the intermediate mass
  • We don't have any of the cabling from the OMC-SUS top to the rack so we can't test anything
  • We haven't uploaded pretty pictures for all to see

We left the OMC partially suspended by the OMC-SUS and partly resting on the installation lab jacks which are currently acting as EQ stops. After we fix the cabling we will more permanently hang it. PS, It looks like the REFL beam extraction will be tricky so we need to get on that....
Attachment 1: IMG_1483.jpg
IMG_1483.jpg
Attachment 2: IMG_1481.jpg
IMG_1481.jpg
  15   Thu Oct 25 22:02:58 2007 robRoutinePSLHEPAs maxed
In light of the SoCal fires, I turned the PSL HEPAs up to 100%.
  14   Thu Oct 25 17:52:45 2007 waldmanOtherOMCOMCs with QPDs
[Rich, Chub, Pinkesh, Sam]

Yesterday we got the QPD, OTAS, and PZT cabling harness integrated with the OMC. We found a few things out, not all of them good. The QPDs went on no problem and could be fairly well aligned by hand. We "aligned" them by looking at all four channels of the QPD on the scope and seeing that there is signal. Since the beam is omega = 0.5 mm, this is a reasonable adjustment. We then connected the OTAS connector to the OTAS and found that the heater on the OTAS was bonded on about 30 degrees rotated from its intended position. This rotated the connector into the beam and caused a visible amount of scattering. This wasn't really a disaster until I removed the connector from the heater and broke the heater off of the aluminum parts of the OTAS. Two steps backwards, one step forward. After the OMC, OMC-SUS integration test we will re-bond the heater to the aluminum using VacSeal. In the meantime, the OMC has been moved to Bridge 056 for integration with the OMC-SUS. More on that as we make progress.
  13   Thu Oct 25 00:01:21 2007 ranaSoftware InstallationCDSGEO DV => LIGO DV
Martin Hewitson of GEO600 fame has modified the cool GEO DV
to work with the LIGO NDS system with some NDS advice from Rolf (who's over in Germany this week).

I've moved it onto the 40m CDS system and installed it on the AdhikariLab computer named 'django'. It worked immediately.

I modified the main .m file to include the 40m's NDS server. When you run it you have to include the path to the NDS
client written by Ben Johnson.

The attached is a screenshot of it working on a Mac; it looks as cool on Linux.

Its installed in /cvs/cds/caltech/apps/ligoDV/. In matlab you navigate to that directory and then
type addpath('/cvs/cds/caltech/apps/linux/UNIX_NDS_Client_beta2/') to add the NDS client.
On the Solaris machines, type type addpath('/cvs/cds/caltech/apps/solaris9/UNIX_NDS_Client_beta2/') instead.

Then type ligoDV to start it up. Then click away and have fun.

In the example I've selected
C1:PEM-BS_ACC_EAST_Z
and plotted its specgram.

Big grin
Attachment 1: Picture_1.png
Picture_1.png
  12   Wed Oct 24 08:58:09 2007 steveOtherPSLlaser headtemp is up
C1:PSL-126MOPA_HTEMP is 19.3C

Half of the chiller's air intake was covered by loose paper
Attachment 1: htempup.jpg
htempup.jpg
  11   Wed Oct 24 01:43:32 2007 Andrey RodionovOtherGeneralPDF-file -> Will report about first results for XARM during Wednesday meeting

Here is the pdf-file with some graphs showing first results for XARM optimization.

We will discuss alltogether during our Wednesday meeting which starts at 2.40PM. Probably it would be necessary to project this pdf-file to the big screen,
so someone should bring laptop and probably connect it to the projector. I do not have a laptop.

See you on that meeting.
Attachment 1: Andrey_October_24.pdf
Andrey_October_24.pdf Andrey_October_24.pdf Andrey_October_24.pdf Andrey_October_24.pdf Andrey_October_24.pdf Andrey_October_24.pdf Andrey_October_24.pdf Andrey_October_24.pdf
  10   Tue Oct 23 11:08:20 2007 steveOtherGeneralbrush fires
There are big brush fires around LA
40 days plot show no effect in the 40m lab
Attachment 1: brushfires.jpg
brushfires.jpg
  9   Tue Oct 23 09:01:00 2007 ranaOtherOMCPZT calibration/ transfer function.
Are you sure that the error signal sweep is not saturated on the top ends? This is usually the downfall
of this calibration method.
  8   Mon Oct 22 19:27:14 2007 pkpOtherOMCPZT calibration/ transfer function.
We measured the PZT transfer function by comparing the PZT response of the circuit with the cavity in the loop, with that of the circuit without the cavity in the loop. Basically measure the transfer function of the whole loop with the laser/PZT and Op-amps in it. Then take another measurement of the transfer function of everything else besides the PZT and from both these functions, we can calculate the PZT response.

The calibration was done by using the error signal response to a triangular wave of volts applied to the PZT. A measurement of the slope of the error signal , which has three zero-crossings as the cavity sweeps through the sidebands, gives us the Volts/Hz response. In order to derive a frequency calibration of the x axis, we assume that the first zero crossing corresponds to the first side band (-29.5 MHz) and the third one corresponds with the other sideband (+29.5 MHz). And then by using the fact that we know the response of the cavity to a constant frequency shift, we can use the Volts/Hz measurement to calculate the Volts/nm calibration. The slope that was calculated was 3.2e-6 V/Hz and using the fact that the cavity is 1 m in length and the frequency is 1064 nm, we get a calibration of 0.9022 V/nm.

Attachment 1: calib.pdf
calib.pdf
Attachment 2: calibpzt2.pdf
calibpzt2.pdf
Attachment 3: all2.pdf
all2.pdf
Attachment 4: noPZT2.pdf
noPZT2.pdf
  7   Mon Oct 22 12:02:59 2007 ajwRoutineGeneralSTACIS as microseismic shaker
In case we ever want to use our Stacis systems as shakers, check this:
link
  6   Sat Oct 20 11:54:13 2007 waldmanOtherOMCOMC and OMC-SUS work
[Rich, Chub, Pinkesh, Chris, Sam]

Friday the 18th was a busy day in OMC land. Both DCPDs were mounted to the glass breadboard and the OMC-SUS structure was rebuilt to the point that an aluminum dummy mass is hanging, unbalanced. The OSEMs have not be put on the table cloth yet, but everything is hanging free. As for the DCPDs, if you recall one beam is 3mm off center from the DCPD tombstone. Fortunately, one DCPD is nearly 3mm offcenter from the case in the right direction, so the errors nearly cancel. The DCPD is too high, so the beam isn't quite centered, but they're close. We'll get photos of the beam positions in someday. Also, the DC gain between the two PDs is, at first glance, different by 15%. DCPD1, the one seen in transmission has 315 mV of signal while DCPD2 has 280 mV. Not sure why, could be because of beam alignment or tolerances in the Preamp or the angle incident on the diode or the QE of the diodes. The glass cans have *not* been removed.
  5   Fri Oct 19 16:11:38 2007 pkpOtherOMCOMC PZT response
Sam and I locked the laser to the OMC cavity and looked at the error signal as a function of the voltage applied to the OMC PZT.
Here are two plots showing the response as a function of frequency from 1 kHz to 100 kHz and another high-res response in the region of 4.5 kHz to 10 kHz.
Attachment 1: fullspec.jpg
fullspec.jpg
Attachment 2: 4.5to10.jpg
4.5to10.jpg
Attachment 3: 4.5to10.pdf
4.5to10.pdf
Attachment 4: fullres.pdf
fullres.pdf
  3   Thu Oct 18 15:03:14 2007 ajwRoutineGeneralthis is only a test

  2   Thu Oct 18 14:52:35 2007 ranaRoutineASCtest
test

X-(:P;(:))
  1   Wed Oct 17 18:46:33 2007 ranaConfigurationGeneraleLog Change
This is the first entry in the new 40m eLog.

Its GWs or bust now! Big grin



[Hnull][/Hnull]
ELOG V3.1.3-